OpenStudy (anonymous):

Solve the equation and check. x-2.9=3.6 Select the correct choice below and, if necessary, fill in the box to complete your choice. A. x=? B. The solution is all real numbers. C. There is no solution.

jimthompson5910 (jim_thompson5910):

you have to move that " -2.9 " over to do that, add 2.9 to both sides

jimthompson5910 (jim_thompson5910):

notice how addition is the opposite of subtraction which allows us to cancel out the "2.9" to isolate x

OpenStudy (xxdragoonxx):

x-2.9=3.6 2.9=29/10,3.6=18/5 = x-29/10=18/5 = x10-29/1010=18/510 = 10x-29=36 = 10x-29+29=36+29 = 10x=65 = 10x/10=65/10 = x=13/2

OpenStudy (anonymous):

@jim_thompson5910 .7?

jimthompson5910 (jim_thompson5910):

we have this originally